OpenStudy (diamond123):

how do yu change a improper fraction into a mixed number then how how do you change it back

OpenStudy (anonymous):

by example \[\frac{17}{5}\] five goes in to 17 three times, remainder is 2 so as a mixed number it is \[3\tfrac{2}{5}\]

OpenStudy (anonymous):

i.e. divide. the whole number is the whole number part and the remainder gets put over the divisor.

OpenStudy (anonymous):

changing back is easy by example \[7\tfrac{3}{5}=\frac{5\times 7+3}{5}=\frac{38}{5}\]

OpenStudy (anonymous):

up top goes denominator times whole number plus numerator. in the bottom goes denominator